3. Along those same lines now we move from moles of ions to mole of charge. Keep in mind that we talk about cmoles instead of moles in soils and 1 cmol = 1/100 of a mole.

show your work!

Ex. 40g Ca2+ x 1mole Ca2+/40 g Ca2+ x 2 molc/1 mole Ca2+=2 molc

1mole H+ = ___________ mol (+)

1mole Ca2+ = ___________ mol (+)

1mole Al3+ = ___________ mol (+)

60g of Ca2+ = ___________ mol (+)

10 g Al3+ = ___________ mol (+)

39 g K+ = _________ cmol (+)
Answer #1

1) 1mole H+ = 1mole H+ x 1mol(+)/1mole H+ = 1mol (+)

2)1mole Ca2+=1mol Ca2+ x 2mol(+)/1mol Ca2+ = 2mol (+)

3) 1mole Al3+ = 1mol Al3+ x 3 mol(+)/1mol Al3+ = 3mol (+)

4) 60g of Ca2+ = 60g of Ca2+ x 1mol Ca2+/40 g Ca2+ x 2mol(+)/1mol Ca2+

= 3mol (+)

5) 10 g Al3+ = 10 g Al3+ x 1mol Al3+/27g Al3+ x 3mol(+)/1mol Al3+

= 1.11mol (+)

6) 39 g K+ = 39 g K+ x 1mol K+/39 g of K+ x 1mol(+)/1mol K+ x 100cmol(+)/1mol(+)

= 100cmol(+)

Please give good rating.